Discuss the intensity of transmitted light when a polaroid sheet is rotated between two crossed polaroids?

Grade:11

1 Answers

Annesha Majumdar
7 years ago
Intensity of light after passing through P2 is=Iocos^2xWhere x is the angle between P1 and P2.Since P1 and P3 are crossed,angle between P2 and P3 is pi/2-xSo,intensity=Iocos^2xcos^2(pi/2-x)=Io/4sin^2(2x)Therefore the intensity is maximum when x=pi/4